OpenStudy (anonymous):

Find all real solutions of the equation. (If there is no real solution, enter NO REAL SOLUTION.) w2 = 8(w − 5) What is: w =____ (smaller value) w =____ (larger value

OpenStudy (anonymous):

is the equation \[w^2=8(w-5)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

\[w^2-8w+40=0 \] thus \[(8\pm \sqrt{64-160})/2\] gives your answer

OpenStudy (anonymous):

which is a complex number

OpenStudy (anonymous):

it has to be 2 answers

OpenStudy (anonymous):

that gives two solutions, but they are not real, so it is no real solution

OpenStudy (anonymous):

I agree. No real solutions